### v2

@@ 4.2 Constitutive models for dilute emulsions
In weak flows free of vorticity, the stationary shape and surfactant distribution are independent of viscosity ratio since Marangoni stresses immobilize the droplet interface [108, 72]. The rheological material functions for drops covered with insoluble surfactants in shear flow are
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5}{2}-D_{2}(\lambda,\beta)Ca^{2}+O(Ca^{3})\,,$ (38)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=\frac{5}{2}\frac{4\beta+1}{\beta}Ca\,,\quad\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{1}{2}\frac{N^{p}_{1}}{\mu\dot{\gamma}}+\frac{75}{28}Ca\,,$ (39)
where the coefficient $D_{2}$ is defined in Appendix A.1 Note that, in the limit of $Ca\to 0$, inserting Eq. (38) into Eq. (23) yields Einstein’s classical result $1+(5/2)\phi$ and emulsion rheology follows the behavior of a suspension of rigid spheres with vanishing normal stress differences.
@@
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5\lambda+2}{2(\lambda+1)}\,,$ (42)
@@
where the small parameter $\epsilon=\lambda^{-1}$ or $Bq_{s}^{-1}$ for $Bq_{s}\sim Bq_{d}$. The form of the coefficients $\hat{a}_{D}$ and $\hat{a}_{E}$ are shown in the Appendix A.2. In this limit, small-deformation theory indicates that the emulsion of either highly viscous internal or surface viscosities behave approximately as rigid spheres with no shear-thinning and no significant elastic effects. This observation is in agreement with the small-deformation theory for high-viscosity drops in weak flows [72, 74].
Droplets with slip at interfaces Ramanchandran & Leal [54] developed a second order small deformation analysis for drops with interfacial slip in weak flows. The model captures the anomalous decrease in relative viscosity measured in emulsions formed by immiscible polymer blends. The viscometric functions in shear flow are
@@
where the functions $f$ and $g$ are defined in Appendix A.3, for completeness [54]. In extensional uniaxial flow, the theory predicts
@@ 5 Non-dilute emulsions: constitutive models and numerical methods
5.1 Constitutive models for semi-dilute and concentrated emulsions
Constitutive equations proposed for non-dilute emulsions aim to account for finite effects of drop deformations, interactions, and microstructure with respect to each other at disperse-phase volume fractions typically above $10\%$. Oldroyd [7] used an effective-medium approach to derive an expression for the effective viscosity of semi-dilute emulsions following a perturbation analysis proposed by Frölich & Sack [65] for suspensions of elastic spheres. Finite effects of disperse-phase volume fraction were included using a cell model. The cell model represents a composite system consisting of a particle surrounded by a volume of suspending fluid, beyond which the emulsion (or a suspension) is seen as a continuum material. This condition is enforced by a modified far-field velocity boundary condition for the disturbance flow generated by the particle (that encompass different particle types, e.g., drops, capsules, vesicles, rigid particles, and blood cells). Specifically, Eq. (4) is evaluated at a truncated far field position $b/a\sim\phi^{-1/3}$, where $b$ is the characteristic size of the cell in which pressure and velocity disturbances are evaluated, and $a$ is particle size. Oldroyd’s effective medium analysis results in
$\eta_{r}=1+\phi\frac{5\lambda+2}{2(\lambda+1)}\left(1+\phi\frac{\left(5\lambda+2\right)}{5(\lambda+1)}\right)\,,$ (50)
for the effective viscosity.
Choi & Schowalter [10] proposed an alternative derivation of effective viscosity of nondilute emulsion by expanding on the stress-averaged, small-deformation theories of Frankel & Acrivos [9] and Cox [8], by accounting for interparticle interactions and higher-order effects of disperse-phase volume fraction. In steady shear flow, the Choi & Schowalter’s constitutive equation yields,
@@ 5.1 Constitutive models for semi-dilute and concentrated emulsions
for relative viscosity of emulsions. The expressions for the normal stress differences are given by Eqs. (30) and (31) in Ref. [10]. The agreement between Eq. (51) and experimental data in Fig. 4 shows that the model up is valid up to terms $O(\phi^{2})$ in the semi-dilute regime. Yaron & Gal-Or [125] proposed a similar model considering a free-surface cell approach to account for surfactant effects in the limit of spherical droplets. Later generalizations of Oldroyd and Choi & Schowalter viscosity models were developed to include non-Newtonian effects of the drop and suspending phases [126, 4], though a lot of open questions remain regarding the influence of viscoelasticity of either liquid or the interface.
@@ 5.2 Empirical equations
$\eta_{r}\left[\frac{2\eta_{r}+5\lambda}{2+5\lambda}\right]=\left(1-\phi/\phi_{m}\right)^{-2.5\phi_{m}}\,.$ (53)
@@
Predictions for Eq. (53) compared to experimental data are shown in Fig. 4. The inset shows data plotted on a linear-linear axis. The corresponding plot shown using log-log scale helps to emphasize how well Taylor’s pioneering theory [6] captures the rheology of dilute emulsions (details about properties of dispersed and suspending liquids are included in Appendix B). The comparison of theory and experiments reveals that the Choi-Schowalter model [10, 4] captures the non-linearity introduced by drop-drop interactions in nondilute emulsions, but the impact of higher order interactions and microstructure require a careful consideration for $\phi$ > 0.4 or so. For a comprehensive review on the empirical viscosity models for concentrated emulsions see Ref. [128]. For nondilute emulsions, normal stress differences become important and shear-thinning effects are also observed at higher shear rates [20]. We recommend a few recent comprehensive reviews for detailed discussion of microstructure, interactions and rheology of concentrated emulsions. [11, 20, 1]
Figure 4: Comparison of theory and empirical relations for effective viscosity models of dilute and concentrated emulsions, respectively. Taylor’s effective viscosity relation is obtained by inserting Eq. (31) into Eq. (23) (dash-dotted line), Choi & Schowalter is given by Eq. (51) (dashed line), and Eq. (53) is used for the Krieger-Dougherty-like curve (dotted line). See Appendix B for details on the datasets and models used.
@@ 5.3 Numerical methods for concentrated emulsions
In this section, we enumerate representative numerical works on modeling semi-dilute to concentrated emulsion flows. We focus on the flow-induced microstructue of deformable drops in unbounded flows. Beyond the dilute regime, pairwise droplet interactions are affected by finite deformation of the drop interface allowing for hydrodynamic diffusion. Droplet deformation in the near contact is the stabilizing mechanism against coalescence in the absence of van der Waals attraction [129]. Scaling analysis for the near-contact motion between two deformable, clean drops within lubrication regime shows a slow algebraic film drainage $h/h_{0}\sim\lambda/(\dot{\gamma}t)$ for $\dot{\gamma}t=O(1)$, where $h$ is the gap between the drops, and $h_{0}$ is a reference, initial gap width. At long times, the internal circulation immobilizes the near-contact motion preventing coalescence [87].
@@
As the disperse-phase volume fraction increases, as illustrated in Fig. 1, many drop interactions become important and an analytical treatment is limited. In this regime, detailed numerical simulations are often used to investigate flow-induced structuring and rheology of concentrated emulsions. The choice of numerical method depends largely on the system parameters (e.g., drop relaxation time, size distribution, and dispersed-phase concentration) and imposed flow conditions. Depending on the type of problem under investigation, for example, whether changes in drop topology or the near contact approach of droplet pairs are of interest, a balance among accuracy, resolution, meshing techniques, and computational cost play a key role in selecting the appropriate numerical method. Complex fluid flows are inherently multiphysics problems governed by phenomena across lengthscales (e.g., from atomistic to continuum descriptions). Continuum numerical approaches for multiphase flows are typically divided in two main categories: interface capturing and interface tracking methods [130, 131].
@@
Interface capturing and tracking methods Interface tracking methods explicitly track marker points on a grid or a mesh that fits the particle interface; classical examples are Boundary Integral Method (BIM) [132] and Immersed Boundary Method (IBM) [133]. Alternatively, interface capturing methods (e.g., Volume of Fluid Method (VoF) [134], Phase Field Method (PFM) [135], and Level Set Method (LSM) [136]) evolve a field variable across the computation domain where the interface is captured implicitly by a specific value of a field variable, for example, the contour of zeroes of the level set function. At continuum scales, where volume-averaged material properties of the fluid are uniform, the interface between two immiscible fluids is often assumed to have zero thickness hence the definition of sharp or dividing interfaces [137]. Interface tracking methods are efficient and accurate in modeling sharp interfaces and are usually the method of choice when physical parameters vary strongly across an interface. However, topological changes (e.g., coalescence and breakup) are challenging and require highly detailed meshing schemes. Interface capturing methods handle topological changes naturally, whereas, interface tracking methods require additional numerical effort. The challenge of using interface capturing methods to model physical systems where material properties are discontinuous across an interface, may be overcome by a hybrid approach of interface capturing methods and immersed interface methods or ghost fluid methods [138, 139, 101].
@@
Appendix C includes as Figure 7 a descriptive map of representative interface tracking, interface capturing, and coarse-grained mesoscopic numerical approaches used in modeling multiphase flows. Figure 5 highlights representative numerical results of concentrated to dense emulsions using some of the methods listed in Fig. 7. For a comprehensive review on numerical methods used in modeling interfacial rheology and sharp-interface methods to solve free-surface flows, the reader is directed to Refs. [93] and [137], respectively; and to Ref. [131] for more details other computational methods for multiphase flows.
@@
Examples of numerical works on concentrated emulsions Loewenberg & Hinch [142] used boundary integral simulations and presented one of the first attempts to simulate small-scale numerical analysis of concentrated emulsion flows of clean, deformable drops with dispersed-phase volume fraction $\phi\leq 30\%$. The results showed a strong shear-thinning behavior, with large positive first and negative normal stress differences, where typically $|N1|>|N2|$. This rheological response is illustrated by the microstrucuture anisotropy shown in Fig. 5(a) where droplets are more deformed and aligned with the flow direction (left image), whereas in the vorticity direction the drops are closely packed (righ image). Elongation of the droplets in the flow direction promotes large $N_{1}$ and facilitates the motion of drops past each other. This droplet arrangement reduces the collisional cross-section and local viscous dissipation leading to a shear thinning behavior. A similar system of interacting droplets in concentrated emulsions with $\phi<30\%$ has been investigated including inertial effects on the emulsion rheology and flow-induced drop structure [150]. The authors used Dissipative Particle Dynamics method where droplets are stabilized against coalescence by a strong repulsive force as illustrated in Fig. 5(i); breakup events are not considered.

### v3

@@ 4.2 Constitutive models for dilute emulsions
In weak flows free of vorticity, the stationary shape and surfactant distribution are independent of viscosity ratio since Marangoni stresses immobilize the droplet interface [123, 77]. The rheological material functions for drops covered with insoluble surfactants in shear flow for surface elasticity parameter $\beta=CaMa$ are
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5}{2}-D_{3}(\lambda,\beta)Ca^{2}+O(Ca^{3})\,,$ (46)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=\frac{5}{2}\frac{(4\beta+1)}{\beta}Ca\,,\quad\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{1}{2}\frac{N^{p}_{1}}{\mu\dot{\gamma}}+\frac{75}{28}Ca\,,$ (47)
where the coefficient $D_{3}$ is defined in Appendix A.1 Note that, in the limit of $Ca\to 0$, inserting Eq. (46) into Eq. (25) yields Einstein’s classical result $1+(5/2)\phi$ given by Eq. (27) with $g_{T}(\lambda)=1$ and emulsion rheology follows the behavior of a suspension of rigid spheres with vanishing normal stress differences.
@@
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5}{2}{g_{T}}\,,$ (50)
@@
where the small parameter $\epsilon=\lambda^{-1}$ or $Bq_{s}^{-1}$ for $Bq_{s}\sim Bq_{d}$. The form of the coefficients $\hat{a}_{D}$ and $\hat{a}_{E}$ are shown in the Appendix A.2. In this limit, small-deformation theory indicates that the emulsions of highly viscous internal or surface viscosities behave approximately as rigid spheres with no shear-thinning and no significant elastic effects. This observation is in agreement with the small-deformation theory for high-viscosity drops in weak flows [77, 78].
Droplets with slip at the interfaces Ramanchandran & Leal [59] developed a second-order small deformation analysis for drops with interfacial slip in weak flows. The model captures the anomalous decrease in relative viscosity measured in emulsions formed by immiscible polymer blends. The viscometric functions in shear flow are
@@
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{(5/2)g_{T}+h(\lambda,\bar{\alpha})}{1+h(\lambda,\bar{\alpha})}+O(Ca^{2})\,,$ (55)
@@
where $h(\lambda,\bar{\alpha})=5\lambda\bar{\alpha}/(\lambda+1)$, and the functions $f$ and $g$ are defined in Appendix A.3, for completeness [59]. In extensional uniaxial flow, the theory predicts
@@
$\frac{\tilde{\mu}/\mu-3}{\phi}=\frac{(5/2)g_{T}+h(\lambda,\bar{\alpha})}{1+h(\lambda,\bar{\alpha})}+\frac{g(\lambda,\bar{\alpha})}{4}Ca+O(Ca^{2})\,,$ (58)
@@
In this section on rheology of dilute emulsions, wall effects and interparticle interactions were neglected. The material functions (specific viscosity and normal stresses) for clean as well as complex interfaces show characteristics of non-Newtonian response, including a nonlinear relation between stresses and the rate of strain giving rise to fluid memory. Frankel and Acrivos [15] extended the works by Chaffey & Brenner [58], Schowalter, Chaffey & Brenner [13] and, Cox [14] and proposed a set of constitutive equations that capture the transient effects of droplet deformation induced by an imposed linear time-dependent flow field. Appendix B revisits the Frankel and Acrivos analysis, and retraces the steps and assumptions that lead them to a three-parameter Jeffreys-like constitutive equation for emulsions of the form
$\Sigma_{ij}+\Lambda\frac{\mathcal{D}\Sigma_{ij}}{\mathcal{D}t}=-p\delta_{ij}+\eta\left(\dot{\gamma}_{ij}+\Lambda_{J}\frac{\mathcal{D}\dot{\gamma}_{ij}}{\mathcal{D}t}\right)\,,$ (60)
where we let $\dot{\gamma}_{ij}=2E_{ij}$. The three parameters are viscosity, and two timescales. The material relaxation time is dependent on the shape relaxation time as follows
$\Lambda=\frac{(2\lambda+3)(19\lambda+16)}{40(\lambda+1)}\tau_{\sigma}=c_{1}(\lambda)^{-1}\tau_{\sigma}\,,$ (61)
and a Jeffreys emulsion retardation time can be determined from:
$\Lambda_{J}=\Lambda\left(1-\frac{4}{5}\frac{\Lambda}{\tau_{\sigma}}\frac{\phi}{\eta_{r}(\lambda,\phi)}\right)\,.$ (62)
@@ 5 Non-dilute emulsions: constitutive models and numerical methods
5.1 Constitutive models based on small deformation and effective medium theories
Constitutive equations proposed for non-dilute emulsions aim to account for finite effects of drop deformations, interactions, and microstructure with respect to each other at dispersed-phase volume fractions typically above $10\%$.
Oldroyd’s effective medium theory (1953) Oldroyd [12] used an effective medium approach to derive an expression for the effective viscosity of the semi-dilute emulsions following a perturbation analysis proposed by Frölich & Sack [70] for suspensions of elastic spheres. Finite size effects of the higher dispersed-phase volume fraction were included using a cell model. The cell model represents a composite system consisting of a drop (or a particle) surrounded by a volume of suspending fluid in a cell beyond which the emulsion (or a suspension) is seen as a continuum material. This condition is enforced by a modified far-field velocity boundary condition for the disturbance flow generated by the particle (here the generalized term particle engulfs all different types, e.g., drops, capsules, vesicles, rigid and deformable particles, and blood cells). Specifically, Eq. (4) is evaluated at a truncated far field position $b/a\sim\phi^{-1/3}$, where $b$ is the characteristic size of the cell in which pressure and velocity disturbances are evaluated, and $a$ is the particle size. Oldroyd’s effective medium analysis results in the following expression for the effective relative viscosity of an emulsion:
$\eta_{r}=1+\phi\frac{5\lambda+2}{2(\lambda+1)}\left(1+\phi\frac{\left(5\lambda+2\right)}{5(\lambda+1)}\right)\,.$ (63)
Using Taylor’s factor ${g_{T}}={(\lambda+2/5)}/(\lambda+1)$ and specific viscosity as $\eta_{sp}$=($\eta_{r}$-1), Oldroyd’s relative viscosity relation can be rewritten in an alternative and compact form as
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{\eta_{sp}}{\phi}=\frac{5}{2}{g_{T}}\left(1+\phi{g_{T}}\right)\,.$ (64)
Choi & Schowalter’s small deformation theory for nondilute emulsions Choi & Schowalter [16] proposed an alternative derivation of effective viscosity of nondilute emulsions by expanding on the stress-averaged, small-deformation theories of Frankel & Acrivos [15] and Cox [14], by accounting for interparticle interactions and higher-order effects of disperse-phase volume fraction. In steady shear flow, Choi & Schowalter’s constitutive equation yields the following expression for relative viscosity of emulsions.
@@ 5.1 Constitutive models based on small deformation and effective medium theories
On comparing their expressions with Oldroyd’s model from 1953, Choi and Schowalter noted that their coefficient for the second order term was 15.6 compared to 2.5 computed by Oldroyd, and the larger coefficient agrees better with the experiments and Oldroyd’s expectation. Alternatively, the shear rheology response can be rewritten in terms of scaled shear stress and normal stress differences as follows:
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{\eta_{sp}}{\phi}=\frac{5}{2}{g_{T}}\left(1+\frac{25}{4}{g_{T}}\phi+O(\phi^{5/3})\right)\,,$ (66)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=\frac{32}{5}{d_{T}}^{2}Ca\phi\,,$ (67)
$\frac{N^{p}_{2}}{N^{p}_{1}}=-\frac{1}{7}\frac{29{\lambda}^{2}+61\lambda+50}{19{\lambda}^{2}+35\lambda+16}.$ (68)
The expressions for the normal stress differences, given by Eqs. (30) and (31) in Ref. [16] are presented here in alternative form. The value of ${N^{p}_{2}}/{N^{p}_{1}}$ goes to, approximately, -0.218 for $\lambda\to\infty$ and -0.446 as $\lambda\to 0$. Furthermore, the Choi-Schowalter model leads to equations for extensional viscosity that reveal a linearly increasing value with strain rate for uniaxial and linearly decreasing value for biaxial extensional flow.
A comparison between the Choi-Schowalter model prediction based on Eq. (65) and experimental data is shown in Fig. 4. The striking agreement emphasizes that the model captures the observations relatively well valid up to terms $O(\phi^{2})$ in the semi-dilute regime. The use of logarithmic x-axis for volume fraction, $\phi$ helps to observe that the original analysis by Taylor [11] captures the linear dependence of $\eta_{r}$ on $\phi$ in the dilute regime reasonably well. Before Choi and Schowalter [16], Yaron & Gal-Or [148] had proposed a similar model considering a free-surface cell approach, that allows for surfactant effects, but without including drop deformation effects. Later generalizations of Frankel & Acrivos, Oldroyd, and Choi & Schowalter viscosity models were developed to include non-Newtonian effects of the drop and suspending phases [149, 4], though a lot of open questions remain regarding the influence of viscoelasticity of the suspending or dispersed liquid phases and the interface.
Figure 4: Comparison of theory and empirical relations for effective viscosity models of dilute and concentrated emulsions, respectively. Taylor’s effective viscosity relation is obtained by inserting Eq. (39) into Eq. (25) (dash-dotted line), Choi & Schowalter is given by Eq. (65) (dashed line), and Eq. (70) is used for the Krieger-Dougherty-like curve (dotted lines) where two dotted lines are added: one using data from Ref. [150] (black), and another one with $\lambda$ adjusted to best fit the experimental data (blue). See Appendix C for details on the datasets and models used.
@@ 5.2 Empirical equations
$\eta_{r}\left[\frac{2\eta_{r}+5\lambda}{2+5\lambda}\right]^{3/2}=\left(1-\phi/\phi_{m}\right)^{-2.5\phi_{m}}\,.$ (70)
@@
Predictions for Eq. (70) compared to experimental data are shown in Fig. 4. The inset shows data plotted on a linear-linear axis. The corresponding plot shown using log-log scale helps to emphasize how well Taylor’s pioneering theory [11] captures the rheology of dilute emulsions (details about properties of dispersed and suspending liquids are included in Appendix C). The comparison of theory and experiments reveals that the Choi-Schowalter model [16, 4] captures the non-linearity introduced by drop-drop interactions in nondilute emulsions, but the impact of higher order interactions and microstructure require a careful consideration for $\phi$ > 0.4 or so. For a comprehensive review on the empirical viscosity models for concentrated emulsions see Ref. [151]. For nondilute emulsions, normal stress differences become important and shear-thinning effects are also observed at higher shear rates[8, 23, 1].
5.3 Doi-Ohta, Maffettone-Minale and other alternatives to small deformation theory
Constitutive models described so far in this article primarily focus on understanding how drop deformation influences and determines emulsion rheology, starting with Eq. (20). The small deformation theory works best for vanishingly low $Ca$ values or in flows with drops closed to spherical. The results of the small deformation theories are recovered by a series of phenomenological models, inspired by the Maffettone-Minale model [152] as briefly reviewed here, that start by assuming that drops are ellipsoidal and allow for drop rotation by vorticity, deformation by strain rate and relaxation to unperturbed state governed by interfacial tension. Slender body theories are used for capturing drop deformation and emulsion rheology for cases when drops are already extended to thread-like shape, usually under high extensional flow, and though transverse cross-section is assumed to remain circular due to surface tension, the drops in forms of long fluid threads stretch and deform along their axial direction [153, 141, 154, 155, 109]. The alternative models are needed as emulsions formed by mechanical mixing of immiscible liquids or by phase separation can contain drops of various sizes and shapes, and during flow, these drops can deform, coalesce and breakup leading to complex evolution of drop size and shape distribution. In 1991, Doi and Ohta [156] proposed an alternative model for emulsion rheology that characterized the rheology in terms of interfacial orientation and area, and three parameters (viscosity, surface tension and volume fraction) but without introducing drop size as an intrinsic length scale. In 1998, Maffettone and Minale [152] provided an alternative phenomenological model for emulsion rheology by assuming ellipsoidal drops and for small to moderate deformation rates. Several variants of ellipsoidal models have been developed since, including for viscoelastic drops or suspensing fluid, and likewise the Doi-Ohta formalism has been extended to allow for variable viscosity ratios, as summarized in several reviews [9, 4, 45, 157, 115].
The Doi-Ohta model and its descendants The Doi-Ohta model, originally proposed for equiviscous and equidense concentrated emulsions [156] adopted a coarse-grained approach to incorporate the influence of change in area and orientation of the interface to the emulsion stress. The Doi-Ohta model describes the total macroscopic stress for an emulsion as:
$\mathbf{\Sigma}=\mathbf{\Sigma}^{0}-\sigma\mathbf{q}\ =2\mu\langle\mathbf{E}\rangle-\langle p\rangle\mathbf{I}-\sigma\mathbf{q}\,,$ (71)
[…]
Here the interface tensor $\mathbf{q}$, a symmetric and second rank tensor, tracks the orientation of the interface where $dS$ is a differential element of interfacial area, V is the system volume, and the integral is over all interfaces. A scalar $Q$ quantifies the amount of total interfacial area per unit volume, and has dimensions of inverse length or 1/length. The interface tensor $\mathbf{q}$ and the parameter $Q$ are related by a pair of evolution differential equations [4, 156]. The original Doi-Ohta model was restricted to a concentrated emulsion, obtained from a 50-50 blend of viscosity-matched and density-matched mixture and incorporated the influence of coarsening on interfacial area and orientation on Q and $\mathbf{q}$ via scaling laws. Subsequent models retain the coarse-grained approach of the Doi-Ohta model, but modify how volume fraction, formation, coalescence, and pinching of drops and coarsening are included and thereafter influence interfacial area and orientation during flow,[156, 9, 4]. A review by Minale [9] provides a concise introduction to the different models inspired by Doi-Ohta formalism.
Ellipsoidal drops based Maffettone-Minale model and its descendants The phenomenological Maffettone and Minale model [152] that assumes that drops are always ellipsoidal and incompressible (drop volume is preserved). The MM model provides the evolution equation for a symmetric, positive-definite, second rank tensor $\mathbf{S}$ with with eigenvalues representing the square semiaxes of the ellipsoid. The tensor $\mathbf{S}$ measures deviations of the droplet from the spherical shape. The non-dimensional version of the equation obtained for scaled shape tensor $\bar{\mathbf{S}}=\mathbf{S}/{a}^{2}$ is as follows.
$\frac{d\bar{\mathbf{S}}}{d\bar{t}}-Ca(\bar{\mathbf{W}}\cdot\bar{\mathbf{S}}-\bar{\mathbf{S}}\cdot\bar{\mathbf{W}})=-{f_{1}}^{MM}[\bar{\mathbf{S}}-g(\bar{\mathbf{S}})\mathbf{I}]$
$+{f_{2}}^{MM}Ca(\bar{\mathbf{E}}\cdot\bar{\mathbf{S}}+\bar{\mathbf{S}}\cdot\bar{\mathbf{E}})\,.$ (73)
The left hand side of the MM equation involve Jaumann derivative rotating with vorticity and the right hand side has the influence of interfacial tension that drives recovery to the unperturbed shape (first term) and viscous drag that deforms the drops (second term). The MM model captures experimental results for simple shear and uniaxial and planar extensional flows as long as drops remain ellipsoidal (limited in Ca range). At O$({Ca})$, the MM model results recover Taylor’s results for the following choices of the two coefficients:
${f_{1}}^{MM}=c_{1}(\lambda),{f_{2}}^{MM}=c_{0}(\lambda)\,.$ (74)
Minale’s review [9] lists and recaps a family of models that were developed inspired by the MM model. The descendant MM models incorporate the influence of coalescence and breakup to capture the response to high deformation rates with some success, and a few variants account for cases when one of the phases is non-Newtonian [4, 9, 152, 157]. Models from MM family that use ellipsoid drops and allow for non-Newtonian drop or matrix phase as well as a few based on small deformation theory are able to capture the linear viscoelastic response of emulsions with reasonable success, but analysis and characterization of nonlinear viscoelasticity and response at high deformation rates remains challenging [157, 149, 4, 9, 115].
@@ 5.4 Numerical methods for concentrated emulsions
In this section, we enumerate representative numerical works on modeling semi-dilute to concentrated emulsion flows. We focus on the flow-induced microstructue of deformable drops in unbounded flows. Beyond the dilute regime, pairwise droplet interactions are affected by finite deformation of the drop interface allowing for hydrodynamic diffusion. Droplet deformation in the near contact is the stabilizing mechanism against coalescence in the absence of van der Waals attraction [158]. Scaling analysis for the near-contact motion between two clean drops within the lubrication regime shows slow algebraic film drainage $h/h_{0}\sim\lambda/(\dot{\gamma}t)$ for $\dot{\gamma}t=O(1)$, where $h$ is the gap between the drops and $h_{0}$ is a reference, initial gap width. At long times, the internal circulation immobilizes near-contact motion preventing coalescence [97].
@@
As the volume fraction of the disperse phase or $\phi$ increases, as illustrated in Fig. 1, many drop interactions become important and analytical treatment is limited. In this regime, detailed numerical simulations are often used to investigate flow-induced structuring and rheology of concentrated emulsions. The choice of numerical method depends largely on the system parameters (e.g., drop relaxation time, size distribution, and dispersed-phase concentration) and imposed flow conditions. Depending on the type of problem under investigation, for example, whether changes in drop topology or the near contact approach of droplet pairs are of interest, a balance among accuracy, resolution, meshing techniques, and computational cost plays a key role in selecting the appropriate numerical method. Complex fluid flows are inherently multiphysics problems governed by phenomena across lengthscales (e.g., from atomistic to continuum descriptions). Continuum numerical approaches for multiphase flows are typically divided in two main categories: interface capturing and interface tracking methods [159, 160].
@@
Interface capturing and tracking methods Interface tracking methods explicitly track marker points on a grid or a mesh that fits the particle interface; classical examples are Boundary Integral Method (BIM) [161] and Immersed Boundary Method (IBM) [162]. Alternatively, interface capturing methods (e.g., Volume of Fluid Method (VoF) [163], Phase Field Method (PFM) [164], and Level Set Method (LSM) [165]) evolve a field variable across the computation domain where the interface is captured implicitly by a specific value of a field variable, for example, the contour of zeroes of the level set function [166]. At continuum scales, where volume-averaged material properties of the fluid are uniform, the interface between two immiscible fluids is often assumed to have zero thickness hence the definition of sharp or dividing interfaces [167]. Interface tracking methods are efficient and accurate in modeling sharp interfaces and are usually the method of choice when physical parameters vary strongly across an interface. However, topological changes (e.g., coalescence and breakup) are challenging and require highly detailed meshing schemes. Interface capturing methods handle topological changes naturally, whereas, interface tracking methods require additional numerical effort. For example, a numerical scheme called Arbitrary Lagrangian-Eulerian (ALE) has been proposed to track moving interfaces and resolve locally singular flows such as the pinching dynamics of fluid droplets [167]. Usually, ALE approach is combined to a spatial and time discretization schemes (e.g., Garlerkin Finite Element (GFEM) and Finite Differences (FDM)) and to specialized meshing techniques to resolve the rapid variations of the pressure and axial velocity in the neck region during pinching [168, 169, 167]. The challenge of using interface capturing methods to model physical systems where material properties are discontinuous across an interface, may be overcome by a hybrid approach of interface capturing methods and immersed interface methods or ghost fluid methods [170, 171, 114].
@@
Figure 5 highlights representative numerical results of concentrated to dense emulsions using some of the methods listed in Fig. 6. Figure 6 compiles a descriptive map of representative interface tracking, interface capturing, and coarse-grained mesoscopic numerical approaches used in modeling of drop deformation in multiphase flows. The first column under each method lists the range of applicability and main features of each method, the second and third columns summarizes their strengths and weaknesses, respectively. Simulating drop deformation in flowing emulsions depend highly on system parameters and on the questions under investigation. For example, if one is interested in evolving the microstructure of an emulsion at low-Reynolds number conditions where topological changes are not relevant to the system dynamics, then using BIM would be a method of choice in terms of accuracy and efficiency compared to interface capturing methods such as VoF or LSM or even coarse-grained methods like LBM. Alternatively, VoF or LSM are more efficient and stable methods compared to BIM to resolve creeping flows of concentrated emulsions where changes in topology are present since no remeshing is needed to handle coalescence or breakup events. However, handling near-contact regions between droplets hinders accuracy and increases the computational cost when using LSM and VoF. Here, our goal is to highlight that each emulsion flow problem introduces inherent physical challenges and should be probed by the appropriate numerical tool. The summary of methods listed in Fig. 6 could be used as a guide.
@@
For a comprehensive review on numerical methods used in modeling interfacial rheology including viscoelastic effects at the interface and sharp-interface methods to solve free-surface flows, the reader is directed to Refs. [106] and [167], respectively; and to Ref. [160] for more details on other computational methods for multiphase flows.
Examples of numerical works on concentrated emulsions Loewenberg & Hinch [174] used boundary integral simulations and presented one of the first attempts to simulate small-scale numerical analysis of concentrated emulsion flows of clean, deformable drops with dispersed-phase volume fraction $\phi\leq 30\%$. The results showed a strong shear-thinning behavior, with large positive first and negative second normal stress differences, where typically $|N_{1}|>|N_{2}|$. This rheological response is illustrated by the microstrucuture anisotropy shown in Fig. 5(a) where droplets are more deformed and aligned with the flow direction (left image), whereas in the vorticity direction the drops are closely packed (right image). Elongation of the droplets in the flow direction promotes large $N_{1}$ and facilitates the motion of drops past each other. This droplet arrangement reduces the collisional cross-section and local viscous dissipation leading to a shear thinning behavior. A similar system of interacting droplets in concentrated emulsions with $\phi<30\%$ has been investigated including inertial effects on the emulsion rheology and flow-induced drop structure [182]. The authors used Dissipative Particle Dynamics method where droplets are stabilized against coalescence by a strong repulsive force as illustrated in Fig. 5(i); breakup events are not considered.